OpenStudy (anonymous):

which of the following is the base-five representation of the number 41 base 4? I know the answer is 32 base 5 but how do you solve it? I don't get it, help please!

OpenStudy (anonymous):

(41) base 4 is 4 times 4 added to 1 times 1 converting the above number to decimal number system we get 4 times 4 is 16 and 1 times 1 is 1 together added to be 17 repeatedly divide the number 17 until you get a remainder less than 5

OpenStudy (anonymous):

"Base" is essentially telling you what the 10's place means. In base 10, 10 is just 10. However in base 4, the digits 10 represent the number 4. So 41 base 4 = 4*4 + 1 = 17 base 10 = 1*10 + 7. You can see that this is the same as 32 base 5 = 3*5 + 2 = 17 base 10. There are different ways to convert bases. You can see that 17/4 = 4 with remainder 1, or 41 base 4. 17/5 = 3 with remainder 2, or 32 base 5.

OpenStudy (anonymous):

oh ok I think I'm getting it

OpenStudy (anonymous):

so what happens if you have 120 base 8?

OpenStudy (anonymous):

100 base 10 = 10*10 = 10^2. Likewise, 100 base 8 = 8*8 = 8^2 = 64. 20 base 8 = 2*8 = 16 0 base 8 = 0. So: 120 base 8 = 1*8^2 + 2*8^1 + 0*8^0 = 64 + 16 + 0 = 80 base 10. In particular, notice how you can write out a number as a sum of integer powers of the base, where the digits are coefficients.

OpenStudy (anonymous):

oh ok I think I get it, so 120 base 8 will equal 80 oh I get its not that hard! thank you! :)